OpenStudy (anonymous):

44% of US adults 18-34 years think that "Made In America" ads boost sales. Another poll show that 78% of adults in the 18-34 years category use social media online. Suppose that 85% of US adults in the 18-34 years category think that "Made in America" ads boost sales or use social media online. If US adult in the 18-34 years category is randomly selected A: What is the probability that the person thinks that " Made in America" ads boost sales and uses social media online? B: What is the probability that made in America ads boost sales given that the person does not use social media onl

OpenStudy (anonymous):

\[85=44+78-x\] solve for \(x\) to find the intersection

OpenStudy (anonymous):

i am using \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] you know the first 3 numbers and aer looking for the fourth

OpenStudy (anonymous):

I dont know were to begin

OpenStudy (anonymous):

ok do you understand the second post i wrote? \[P(A\cup B)=P(A)+P(B)-P(A\cap B)\]

OpenStudy (anonymous):

the number on the right side is the 85% that is the total of one or the other ( or both) if A is the people who agree with made in america, that is 44% and if B is the social media people, that is 78% now \(85=44+78-P(A\cap B)\) or \[85=122-P(A\cap B)\]so \(P(A\cap B)=122-85=37\)

OpenStudy (anonymous):

this is the answer to the first question, the number of people in both categories is \(37\%\)

OpenStudy (anonymous):

now it would probably be easiest to draw a venn diagram. are you familiar with those?

OpenStudy (anonymous):

no